OpenStudy (help108):

1 math question MEDAL AND FAN!!!1 Consider the following set of equations: Equation A: y = −x + 5 Equation B: y = 6x − 2 Which of the following is a step that can be used to find the solution to the set of equations? −x = 6x + 2 −x − 2 = 6x + 5 −x + 5 = 6x – 2 −x + 5 = 5x

OpenStudy (help108):

@AloneS

OpenStudy (help108):

@iGreen

alones (alones):

Igreen left OS already

OpenStudy (help108):

i know lol

OpenStudy (help108):

but why?

OpenStudy (shadowlegendx):

What do you think the answer is?

OpenStudy (help108):

C someone taught me and said it was C i understood it a little but they werent sure because to find the answer 6-2 and for the other one you do 5+x

OpenStudy (shadowlegendx):

You're correct, it is C. But you want to know why it's correct?

OpenStudy (help108):

thanks mate!!!!

OpenStudy (shadowlegendx):

Basically, when you are trying to solve for variable(s), you want to isolate them. Equation A: y = −x + 5 Equation B: y = 6x − 2 You can't solve for x or y, with both of them being there, we need to get rid of one of them We see that y = -x + 5 , and y = 6x - 2 This means that -x + 5 = 6x - 2 is a true equation

OpenStudy (help108):

okay

OpenStudy (shadowlegendx):

From there you can solve. - x + 5 = 6x - 2 Add x to both sides 5 = 7x - 2 Add 2 to both sides 7 = 7x Divide both sides by 7 1 = x x = 1

OpenStudy (shadowlegendx):

Input back in -(1) + 5 = 6(1) - 2 -1 + 5 = 6 - 2 4 = 4

OpenStudy (shadowlegendx):

x = 1